# TaSL: Task Skill Localization and Consolidation for Language Model Continual Learning

### v1

@@ Abstract
Language model continual learning (CL) has recently garnered significant interest due to its potential to adapt large language models (LLMs) to dynamic real-world environments without re-training. A key challenge in this field is catastrophic forgetting, where models lose previously acquired knowledge when learning new tasks. Existing methods commonly employ multiple parameter-efficient fine-tuning (PEFT) blocks to acquire task-specific knowledge for each task, but these approaches lack efficiency and overlook the potential for knowledge transfer through task interaction. In this paper, we present a novel CL framework for language models called Task Skill Localization and Consolidation (TaSL), which enhances knowledge transfer without relying on memory replay. TaSL first divides the model into ‘skill units’ based on parameter dependencies, enabling more granular control. It then employs a novel group-wise skill localization technique to identify the importance distribution of skill units for a new task. By comparing this importance distribution with those from previous tasks, we implement a fine-grained skill consolidation strategy that retains task-specific knowledge, thereby preventing forgetting, and updates task-shared knowledge, which facilitates bi-directional knowledge transfer. As a result, TaSL achieves a superior balance between retaining previous knowledge and excelling in new tasks. TaSL also shows strong generalizability, suitable for general models and customizable for PEFT methods like LoRA. Additionally, it demonstrates notable extensibility, allowing integration with memory replay to further enhance performance. Extensive experiments on two CL benchmarks, with varying model sizes (from 220M to 7B), demonstrate the effectiveness of TaSL and its variants across different settings.
@@
Language model continual learning, catastrophic forgetting, skill localization.
@@ I Introduction
Equipping large language models (LLMs) with continual learning (CL) capabilities to sequentially learn different tasks is essential for their deployment in real-world scenarios [1, 2]. This capability enables LLMs to dynamically adapt to new tasks and acquire additional knowledge [3, 4]. An effective CL system must address two critical challenges: (1) Catastrophic Forgetting (CF)  [5], the phenomenon where a model’s proficiency in previous tasks deteriorates as it learns new ones, and (2) Knowledge Transfer (KT)  [6], enhancing task performance through the transfer of knowledge. KT can be categorized into forward transfer, which improves new task performance using knowledge from previous tasks, and backward transfer, which enhances performance on previous tasks after learning a new relevant task. Achieving a balance between retaining previous knowledge and excelling in new tasks is vital for success.
@@
Figure 1: Conceptual illustration of TaSL. By identifying task-relevant areas across both previously accumulated and current tasks, we can consolidate the task-specific and task-shared parameters to facilitate efficient knowledge transfer (KT) and mitigate catastrophic forgetting (CF).
@@
Given the considerable computational demands, recent efforts have explored CL for LLMs using parameter-efficient fine-tuning (PEFT) methods [7, 8], such as Low-Rank Adaptation (LoRA) [7]. Traditional rehearsal-based approaches, which require storing data examples from past tasks for replay, face significant privacy and memory issues [9, 10, 3].
@@
Another prominent line of work in recent studies is the parameter isolation CL methods [11, 12, 13, 14]. These methods assign a dedicated PEFT block for each new task to acquire task-specific knowledge, preserving the block for selective activation during testing. However, they exhibit significant limitations in effectively addressing the KT and CF challenges.
@@
On one hand, the prevalent design of these frameworks, which utilizes multiple PEFT blocks, introduces redundancy and inefficiency [15]. The number of required PEFT blocks increases with each new task, leading to substantial memory storage costs and making it difficult to handle long task sequences [16, 17]. Moreover, such approaches essentially create separate expert models for each task, which limits their ability to generalize to unseen tasks [18] (Limitation 1).
@@
On the other hand, these parameter-isolation CL methods do not consider interactions between tasks, which hinders KT. For instance, the method in [12] involves learning each PEFT block separately within individual tasks. Similarly, Orthogonal Low-Rank Adaptation (O-LoRA) [19] updates parameters via gradient projection in orthogonal subspaces. Although these approaches may mitigate CF to some extent, they cut off the potential transfer of knowledge stored across different PEFT blocks, thus impeding bi-directional KT among various tasks and leading to suboptimal performance (Limitation 2).
@@
To address these limitations, we introduce Task Skill Localization and Consolidation (TaSL) [20], a novel CL framework designed to improve KT between tasks without relying on memory replay. Our approach is motivated by recent findings that model parameters contribute unevenly to performance [21]. For instance, the authors in [22] uncovered a core region on LLMs crucial for all languages, while distinct monolingual regions exist for different languages. Preserving these important regions can help avoid forgetting. The research in [15] also found that during fine-tuning with LoRA, many redundant parameter modifications are often retained, suggesting that PEFT blocks contain significant redundancies.
@@
Based on these insights, TaSL facilitates KT by identifying and consolidating the importance distribution of model parameters across tasks. TaSL initially employs a group-wise importance-aware skill localization technique that utilizes gradient trajectories to identify tiny regions within the parameter space that store crucial knowledge for the current task. By comparing the importance distribution with those of previous tasks, we can then differentiate between task-specific and task-shared regions, as illustrated in Figure 1. Our innovative skill consolidation phase then categorically integrates weights from previous tasks with the current one, enabling effective KT while minimizing forgetting.
@@
The importance-aware skill localization method employs a new group-wise metric to compute importance scores, effectively quantifying the significance of each skill unit for the current task. Our approach, focusing on parameter space rather than dataset-driven categorization of task-specific and task-shared knowledge, offers a more effective solution to managing KT in LLMs, overcoming inaccuracies caused by dataset noise. Our skill consolidation stage, then based on a fine-grained model averaging strategy, effectively manages different types of knowledge. During this phase, we promote forward KT by initializing new tasks with previously fine-tuned weights. For backward KT, we merge knowledge from both current and past tasks into localized task-shared skill units, enhancing their capability. To prevent CF, we preserve the integrity of skill units containing previous task-specific knowledge, ensuring they remain unaffected by new task learning (addressing Limitation 2).
Given the widespread adoption and success of LoRA in fine-tuning LLMs, there is a compelling opportunity to optimize the TaSL framework specifically for LoRA. While the original TaSL framework has shown promising results, it has only been validated on a specific sub-task, and its effectiveness on general CL benchmarks has yet to be explored. Additionally, on the technical side, the skill localization process in TaSL relies on first-order gradients, which may not always provide precise importance localization. Furthermore, the skill consolidation phase in TaSL involves many hyperparameters, potentially complicating the averaging process.
To address these issues, we propose TasLoRA, a LoRA-tailored version of TaSL. TasLoRA reconstructs the LoRA adapter into new skill units based on parameter dependencies, ensuring more efficient management of knowledge through sequential task learning. During the fine-tuning process, an additional orthogonal loss is added to ensure that skill units within the same LoRA adapter capture distinct latent semantic features. For skill localization, TasLoRA employs a novel second-order gradient approximate group-wise metric to compute importance scores more precisely. For skill consolidation, we transition from the previous hard-mask model averaging approach to a soft-masking strategy, using an adaptive averaging technique. This allows for a more flexible integration of task-specific and shared parameters, dynamically adjusting the contribution of each skill unit based on its importance.
Furthermore, to adapt to a broader range of application scenarios, we combine TaSL with memory replay, introducing the TaSL-M model. Through extensive experiments on two CL benchmarks with different parameter-level backbones (from 220M to 7B), our TaSL framework, along with its variants, excels in mitigating CF and showcases remarkable capabilities for KT, outperforming state-of-the-art (SOTA) methods.
@@
  • We propose a novel Task Skill Localization and Consolidation (TaSL) framework for language model continual learning. By identifying and consolidating task-specific and task-shared knowledge at a granular skill unit level, TaSL achieves effective knowledge transfer and mitigates catastrophic forgetting, overcoming the limitations of previous approaches.
  • We develop various group-wise skill localization and fine-grained skill consolidation techniques. For instance, the parameter importance metric in skill localization can be based on first-order gradients or a new second-order gradient approximation method. In skill consolidation, our model averaging strategies include a categorical hard-mask approach and an adaptive soft-mask method.
  • The TaSL framework demonstrates strong generalizability and extensibility. The flexible design of skill units enables TaSL to be easily tailored to PEFT methods, such as LoRA, optimizing performance for specific model architectures. Additionally, TaSL can be integrated with memory replay technique to further enhance performance and adapt to a broader range of application scenarios.
  • Extensive evaluation on two CL benchmarks underscores the superiority of our TaSL framework and its variants in facilitating knowledge transfer and mitigating catastrophic forgetting, especially in memory-free scenarios. Furthermore, TaSL consistently excels across diverse model sizes (ranging from 220M to 7B), different model architectures (T5 and LLaMA-2), and unseen tasks.
@@ II Related Work
II-A Continual Learning
Continual learning [3] aims to develop learning algorithms capable of accumulating knowledge from non-stationary data.
Conventional Continual Learning methods can be divided into three categories: (i) Regularization-based methods add explicit regularization terms to preserve the knowledge of previous tasks [13, 23]. (ii) Rehearsal-based methods address catastrophic forgetting by storing old training samples [24, 25] or by training generative models to provide pseudo samples of previous tasks [26, 27]. Due to the memory and data privacy issues associated with these methods, we focus on rehearsal-free CL methods. (iii) Parameter isolation-based methods dynamically expand model capacity or assign isolated parameters dedicated to each task throughout the CL process to prevent interference between tasks [28, 29, 30].
Language Model Continual Learning with PEFT. Based on PEFT methods, current approaches for the continual learning of LLMs adopt the concept of parameter isolation, using a pipeline fashion to learn and select PEFT blocks for each task [12, 31, 32]. However, these approaches are limited in effectively addressing the challenges of catastrophic forgetting (CF) and knowledge transfer (KT) [33]. For instance, the authors in [19] and [34] constrain the learning of PEFT blocks to maintain orthogonality, restricting KT among different tasks. Although the recent SAPT method [16] and DAP method [35] achieves KT, it relies on pseudo sample replay or unlabeled domain corpora.
Our method achieves KT without memory replay, offering unique advantages in data privacy and parameter efficiency. By using layering low-rank adapters on the key and value projection matrices of transformer blocks, we balance retaining previous knowledge and excelling in new tasks.
@@ II-B Skill Localization
Research shows that model parameters do not all contribute equally to performance [36]. The authors in [21] introduced the concept of “skill localization” to identify critical parameters in pre-trained language models, suggesting that fine-tuning only these critical parameters can achieve results comparable to full model fine-tuning. However, their method involves additional steps to identify and retrain these key parameters during the post-fine-tuning phase, which reduces efficiency.
Inspired by techniques from the pruning community, previous studies have utilized gradient-based metrics to pinpoint important parameters during fine-tuning. Sensitivity-based scoring [37, 38] evaluates the impact on training loss, and sensitivity smoothing, as employed by [39], helps remove unnecessary parameters for more efficient fine-tuning. Nonetheless, these methods often result in element-wise pruning, where individual parameters are removed based on their importance scores. This approach can be computationally and storage-intensive, as it requires detailed management of numerous parameters throughout the model.
Building on these advancements, we propose a novel importance-aware skill localization method. This approach uniquely differentiates between task-specific and shared parameters to mitigate forgetting and achieve KT in CL.
@@ III Preliminaries
III-A Continual Learning
Continual learning [40] aims to create learning algorithms that can continuously accumulate knowledge from ongoing sequences. In supervised continual learning, a sequence of tasks $\{\mathcal{T}_{1},\ldots,\mathcal{T}_{K}\}$ is presented in a streaming manner. Each task $\mathcal{T}_{k}$ includes a distinct target dataset $\mathcal{D}_{k}=\left\{\left(x_{i}^{k},y_{i}^{k}\right)\right\}_{i=1}^{N_{k}}$ of size $N_{k}$, where $x_{i}^{k}\in\mathcal{X}_{k}$ and $y_{i}^{k}\in\mathcal{Y}_{k}$. The model must adapt to these tasks sequentially, having access only to $\mathcal{D}_{k}$ during the $k$-th task. Generally, with a prediction model $f$ parameterized by $\Theta$, continual learning aims to optimize the following objective across all tasks:
@@ III-A Continual Learning
The notation $f_{k}$ refers to the model after training on task $\mathcal{T}_{k}$, while $\hat{f}_{k}$ denotes the model after averaging for $\hat{f}_{k-1}$ and $f_{k}$. Our TaSL framework aims to address a more challenging scenario where the model cannot access any historical data during training [19]. Nevertheless, to adapt to a wider range of application scenarios, we have also combined TaSL with memory replay. In this setting, we randomly save $\left|\mathcal{M}\right|$ samples from the training set of each previous task $\mathcal{T}_{i}$ in memory $\mathcal{M}_{i}$ and jointly train the model on new task data $\mathcal{D}_{k}$ and memory $\mathcal{M}_{<k}$. This extension, called TaSL-M, leverages memory replay to enhance performance.

### v2

@@ Abstract
Language model continual learning (CL) has recently attracted significant interest for its ability to adapt large language models (LLMs) to dynamic real-world scenarios without retraining. A major challenge in this domain is catastrophic forgetting, where models lose previously acquired knowledge upon learning new tasks. Existing approaches commonly utilize multiple parameter-efficient fine-tuning (PEFT) blocks to acquire task-specific knowledge, yet these methods are inefficient and fail to leverage potential knowledge transfer across tasks. In this paper, we introduce a novel CL framework for language models, named Task Skill Localization and Consolidation (TaSL), which boosts knowledge transfer without depending on memory replay. TaSL initially segregates the model into ‘skill units’ based on parameter dependencies, allowing for more precise control. Subsequently, it employs a novel group-wise skill localization technique to ascertain the importance distribution of skill units for a new task. By comparing this importance distribution with those from previous tasks, we implement a fine-grained skill consolidation strategy that retains task-specific knowledge, thereby preventing forgetting, and updates task-shared knowledge, which facilitates bi-directional knowledge transfer. As a result, TaSL achieves an optimal balance between retaining prior knowledge and excelling in new tasks. TaSL also demonstrates strong generalizability, making it suitable for various base models and adaptable to PEFT methods like LoRA. Furthermore, it offers notable extensibility, supporting enhancements through integration with memory replay techniques. Comprehensive experiments conducted on two CL benchmarks, involving models ranging from 220M to 7B parameters, affirm the effectiveness of TaSL and its variants across different settings.
@@
Language model continual learning, catastrophic forgetting.
@@ I Introduction
Equipping large language models (LLMs) with continual learning (CL) capabilities to sequentially learn different tasks is essential for their deployment in real-world scenarios [1, 2]. This capability enables LLMs to dynamically adapt to new tasks and acquire additional knowledge [3, 4]. An effective CL system must address two critical challenges: (1) Catastrophic Forgetting (CF)  [5], where a model’s performance on previous tasks deteriorates as it learns new ones, and (2) Knowledge Transfer (KT)  [6], enhancing task performance through the transfer of knowledge. KT can be categorized into forward transfer, which improves new task performance using knowledge from previous tasks, and backward transfer, which enhances performance on previous tasks after learning a new relevant task. Achieving a balance between retaining previous knowledge and excelling in new tasks is vital for success.
@@
Figure 1: Conceptual illustration of TaSL. By identifying task-relevant areas across both previously accumulated and current tasks, we can consolidate the task-shared and task-specific parameters to facilitate efficient knowledge transfer (KT) and mitigate catastrophic forgetting (CF).
@@
Due to the high computational demands, recent studies have explored CL for LLMs using parameter-efficient fine-tuning (PEFT) methods [7, 8], such as Low-Rank Adaptation (LoRA) [7]. Traditional rehearsal-based approaches, which require storing data examples from past tasks for replay, face significant privacy and memory issues [9, 3].
@@
Another prominent line of work in recent studies is the parameter isolation CL methods [10, 11, 12, 13]. These methods assign a dedicated PEFT block for each new task to capture task-specific knowledge, preserving the block for selective activation during testing. However, they exhibit significant limitations in effectively addressing the KT and CF challenges.
@@
On one hand, the prevalent design of these frameworks, characterized by the use of multiple PEFT blocks, introduces redundancy and operational inefficiency [14]. As the number of tasks grows, the required PEFT blocks accumulate, leading to increased memory storage demands and making it difficult to handle long task sequences [15, 16]. Furthermore, such approaches tend to develop separate specialist models for each task, which severely limits their capability to generalize across unseen tasks [17] (Limitation 1).
@@
On the other hand, these parameter-isolation CL methods overlook the interactions between tasks, which hinders KT. For instance, the method in [11] learns each PEFT block separately within individual tasks. Similarly, Orthogonal Low-Rank Adaptation (O-LoRA) [18] employs orthogonal subspace gradient projections for parameter updates. While these strategies might mitigate CF, they cut off the potential for leveraging knowledge distributed across various PEFT blocks, thus obstructing bi-directional KT among tasks and resulting in less-than-optimal performance (Limitation 2).
@@
To address these limitations, we introduce Task Skill Localization and Consolidation (TaSL) [19], a novel CL framework designed to improve KT between tasks without relying on memory replay. Our approach is motivated by recent findings that model parameters contribute unevenly to performance [20]. For instance, the authors in [21] uncovered a core region in LLMs that is crucial for all languages, suggesting that preserving these vital regions can mitigate forgetting. Additionally, findings from [14] indicate that fine-tuning with LoRA often preserves many unnecessary parameter changes, pointing to substantial redundancies within PEFT blocks.
@@
Based on these insights, TaSL facilitates KT by pinpointing and consolidating the importance distribution of model parameters across tasks. Initially, TaSL utilizes a group-wise importance-aware skill localization technique that employs gradient trajectories to identify tiny regions within the parameter space that store crucial knowledge for the current task. By comparing the importance distribution with those of previous tasks, we can then differentiate between task-specific and task-shared regions, as illustrated in Figure 1. The subsequent skill consolidation phase then categorically integrates weights from previous tasks with the current one, efficiently mitigating CF.
@@
The importance-aware skill localization method employs a new group-wise metric to compute importance scores, effectively quantifying the significance of each skill unit for the current task. Our skill consolidation phase, then based on a fine-grained model averaging strategy, effectively manages different types of knowledge. This stage facilitates forward KT by using previously fine-tuned weights as the starting point for new tasks. For backward KT, we merge knowledge from both current and past tasks into localized task-shared skill units, boosting their effectiveness. To prevent CF, we ensure the integrity of skill units containing prior task-specific knowledge is maintained, safeguarding them from being altered by the learning of new tasks (addressing Limitation 2).
Given the widespread adoption and success of LoRA in fine-tuning LLMs, there is a compelling opportunity to optimize the TaSL framework specifically for LoRA. While the initial TaSL framework has shown promising results, its validation has been limited to a specific sub-task, and its broader applicability on general CL benchmarks remains explored. Additionally, the reliance on first-order gradients for skill localization in TaSL may not yield precise importance localization. The skill consolidation phase also involves many hyperparameters, which could complicate the averaging process.
To address these issues, we propose TasLoRA, a variant of TaSL optimized for LoRA. TasLoRA redesigns the LoRA adapter into new skill units based on parameter dependencies, streamlining knowledge management throughout sequential task learning. It incorporates an orthogonal loss during fine-tuning to ensure distinct latent semantic features are captured within the same LoRA adapter. For skill localization, TasLoRA utilizes a new second-order gradient approximate group-wise metric, enhancing the precision of importance scores. In skill consolidation, it shifts from hard-mask to soft-mask averaging, employing an adaptive technique that flexibly integrates task-specific and shared parameters according to the importance of the skill unit. Additionally, we combine TaSL with memory replay to broaden its application scope, introducing the TaSL-M model. Through extensive experiments on two CL benchmarks with four parameter-level backbones, our TaSL framework and its variants excel in mitigating CF and showcase remarkable capabilities for KT, outperforming SOTA methods.
@@
  • We introduce the Task Skill Localization and Consolidation (TaSL) framework for language model continual learning. By pinpointing and amalgamating task-specific and task-shared knowledge at a granular skill unit level, TaSL ensures effective knowledge transfer and significantly reduces catastrophic forgetting, addressing the shortcomings of previous methodologies.
  • We devise various group-wise skill localization and fine-grained skill consolidation techniques. Notably, our skill localization utilizes either first-order gradients or an innovative second-order gradient approximation for assessing parameter importance. For skill consolidation, we implement strategies ranging from categorical hard-masking to adaptive soft-masking, enhancing the flexibility and precision of knowledge integration.
  • The TaSL framework is distinguished by its robust generalizability and extensibility. Its skill units are designed flexibly, allowing for seamless tailoring to PEFT methods, such as LoRA, optimizing performance across different model architectures. Moreover, integrating TaSL with memory replay enables further performance enhancements and broadens its applicability to diverse scenarios.
  • Extensive evaluation on two CL benchmarks demonstrates the superiority of our TaSL framework and its variants in facilitating knowledge transfer and mitigating catastrophic forgetting, especially in memory-free scenarios. Furthermore, TaSL consistently excels across diverse model sizes (from 220M to 7B), various model architectures (T5 and LLaMA-2), and unseen tasks.
@@ II Related Work
II-A Language Model Continual Learning
Continual learning [3] seeks to develop algorithms that accumulate knowledge from non-stationary data sources.
Conventional continual learning methods can be categorized into three main types: (i) Regularization-based methods impose explicit constraints to preserve the knowledge of previous tasks by penalizing the variation of each parameter based on its contribution to past tasks [12, 22], such as in the EWC [23] method. However, these methods typically restrict the learning of task-shared knowledge by only considering parameter importance from the perspective of past tasks. This limitation often leads to a suboptimal balance between retaining previous knowledge and excelling at new tasks, as it does not account for the relevance of parameters to both past and current tasks. (ii) Rehearsal-based methods mitigate catastrophic forgetting by either storing old training samples [24, 25] or training generative models to provide pseudo-samples of previous tasks [26, 27], enabling data replay during the learning of new tasks. (iii) Architecture-based methods aim to prevent task interference by dynamically expanding model capacity or assigning dedicated parameter subspaces to each task [28, 29]. While effective at minimizing forgetting, these methods often hinder knowledge transfer (KT) between tasks. Our TaSL framework stands out by bi-directionally analyzing parameter importance across both historical and current tasks, enabling effective KT while mitigating forgetting. To further enhance performance, we extend TaSL with memory replay techniques, resulting in TaSL-M, improving its flexibility and applicability across various scenarios.
Language model continual learning with PEFT methods in the era of LLMs often adopt a parameter isolation strategy. This approach involves using a pipeline to learn and select PEFT blocks for each task [11, 30]. However, as the number of tasks increases, the accumulation of PEFT blocks can become inefficient, and achieving KT between blocks remains challenging. More recently, authors in [31] have explored using the Mixture of Experts (MoE) approach to connect multiple PEFT blocks, but this essentially relies on coarse-grained model averaging. This can lead to overemphasizing unimportant weights, contaminating previously acquired task-specific knowledge, and causing forgetting. To address these challenges in the context of PEFT-based CL, we extend the TaSL framework and introduce TasLoRA. By dividing fine-grained skill units based on parameter dependencies within LoRA, TasLoRA enhances KT efficiency, enables more precise control, and effectively mitigates forgetting.
@@ II-B Skill Localization
Research has shown that model parameters contribute unevenly to performance [32]. For example, the authors in [21] identified a core region in LLMs linked to linguistic competence; freezing this region during further pre-training has been shown to mitigate catastrophic forgetting. Additionally, the authors in [33] found neurons within LLMs that store personally identifiable information, highlighting the potential for improved privacy protections. However, these studies primarily focus on describing such findings rather than applying them to solve practical problems.
Motivated by these findings, we address the challenge of catastrophic forgetting in CL. While [20] introduced the concept of “skill localization” to identify critical parameters in pre-trained language models, their approach requires additional steps to identify and retrain these parameters post-fine-tuning, which impacts efficiency. In contrast, our importance-aware skill localization method leverages trajectory gradients during each task’s training phase to identify parameter importance, significantly reducing processing time. Combined with our skill consolidation strategy, which explicitly targets task-specific and shared parameters, our approach effectively mitigates forgetting and achieves knowledge transfer in CL.
@@ III Preliminaries
III-A Continual Learning Setup
Continual learning [2] aims to develop algorithms that can progressively accumulate knowledge from ongoing sequences. In supervised continual learning, a series of tasks $\{\mathcal{T}_{1},\ldots,\mathcal{T}_{K}\}$ is presented sequentially. Each task $\mathcal{T}_{k}$ includes a distinct target dataset $\mathcal{D}_{k}=\left\{\left(x_{i}^{k},y_{i}^{k}\right)\right\}_{i=1}^{N_{k}}$ of size $N_{k}$, where $x_{i}^{k}\in\mathcal{X}_{k}$ and $y_{i}^{k}\in\mathcal{Y}_{k}$. The model, parameterized by $\Theta$, is trained to adapt to these tasks one at a time, accessing only $\mathcal{D}_{k}$ during the $k$-th task. The overarching goal of continual learning is to optimize the following objective:
@@ III-A Continual Learning Setup
And we denote $f_{k}$ as the model trained up to task $\mathcal{T}_{k}$, while $\hat{f}_{k}$ represents the model after integrating knowledge from $\hat{f}_{k-1}$ and $f_{k}$ through averaging. Our TaSL framework is designed to tackle a more challenging scenario where the model cannot access any historical data during training [18]. To broaden its applicability, we have augmented TaSL with memory replay. In this enhanced setup, known as TaSL-M, we store a random subset of $\left|\mathcal{M}\right|$ samples from each prior task’s training set $\mathcal{D}_{i}$ in memory $\mathcal{M}_{i}$. The model is then trained jointly on the new task data $\mathcal{D}_{k}$ and the accumulated memory $\mathcal{M}_{<k}$, leveraging this memory replay to improve performance.